# ChatCFD: an End-to-End CFD Agent with Domain-specific Structured Thinking

### v1

@@ 3 Results and discussion
The performance of ChatCFD was assessed through validation experiments on basic CFD examples and two complex cases, using published literature as prompts. ChatCFD achieved error-free configurations in 30-40% of cases and operational success in 60-80%, marking a significant advance in automated CFD setup. The system demonstrated robust handling of complex configurations, particularly those with diverse turbulence models, highlighting its adaptability to varied physical scenarios. Analysis revealed two key performance factors: the extent of physical coupling in the flow and the availability of comprehensive knowledge about relevant physical models. These findings elucidate the capabilities and limitations of AI-driven CFD automation, contributing to the advancement of automated CFD workflows.
3.1 CFD cases for Experiments
Figure 3: Demonstration of the two tested CFD cases, including grids of (a) the Naca0012 case [29] and (b) the Nozzle case [36], and boundary types of (c) the Naca0012 case and (d) the Nozzle case.
@@ 3.1 CFD cases for Experiments
We selected two representative test cases for simulation: an incompressible flow and a compressible flow case (illustrated in Figure 3). These cases were chosen based on the comprehensive documentation available in their source articles, which provided detailed specifications of the solver configurations, turbulence models, and mesh characteristics essential for accurate CFD reproduction. For the incompressible flow case, we utilized the work of Sun et al.[29], who conducted simulations of a NACA0012 airfoil at 10° angle of attack employing simpleFoam solver with the SpalartAllmaras turbulence model. For the compressible flow case, we referenced Yu et al.[36], who performed simulations of a nozzle with pressure ratio 3. The rhoCentralFoam and SpalartAllmaras model are used in the Nozzle case.
Given the documented challenges in modifying OpenFOAM turbulence models using LLM, as reported by Pandey et al.[25], we extended our investigation to include additional experiments with the NACA0012 case implementing various turbulence models. Table 1 presents a comprehensive overview of all case configurations and experimental iterations conducted in this study. High-quality computational meshes for these experiments were generated using commercial software. The experimental design featured a structured testing protocol: Cases 1 and 5 were subjected to an extensive model ablation analysis, with 10 iterations performed for each of the five system configurations. In contrast, Cases 2 through 4 were evaluated using only the complete system configuration, also with 10 iterations each. This systematic methodology yielded a total of 130 ChatCFD experimental runs, forming a robust dataset for subsequent analysis.
@@
3.2 Model Ablation Analysis
@@
To systematically evaluate the contributions of different components and strategies, ablation experiments were performed using five distinct system configurations. These configurations vary in terms of comprehensive paper interpretation detailed in Stage 2 (Initiating Case Files), the sophistication of error handling logic, and the utilization of retrieval modules (ReferenceRetriever and ContextRetriever) in Stage 3 (Correcting Case Files):
@@
  • Configuration A (Baseline System): Utilizes a simplified parameter extraction process instead of the three-step process in Stage 2. In Stage 3, error correction is limited to a basic “general error” pathway, with both ReferenceRetriever (tutorial-based retrieval) and ContextRetriever (current case retrieval) modules deactivated.
  • Configuration B (ReferenceRetriever with Basic Error Handling): Maintains the simplified parameter extraction (Stage 2) and basic ’general error’ handling (Stage 3) of Configuration A, but activates the ReferenceRetriever module to inform corrections with tutorial case data.
  • Configuration C (Multi-Category Errors, ReferenceRetriever, Simplified Extraction): Employs the simplified parameter extraction in Stage 2. However, Stage 3 incorporates the full multi-category error correction logic (addressing dimensional, missing file, repetitive, and general errors) and activates the ReferenceRetriever module. The ContextRetriever module remains deactivated.
@@
  • Configuration D (Multi-Category Errors, ReferenceRetriever & ContextRetriever, Simplified Extraction): Also uses the simplified parameter extraction in Stage 2. Stage 3 employs the full multi-category error correction logic, supported by both ReferenceRetriever and ContextRetriever modules.
  • Configuration E (Complete System): Represents the most comprehensive setup. It features comprehensive paper interpretation in Stage 2. In Stage 3, it utilizes the full multi-category error correction logic, augmented by both ReferenceRetriever and ContextRetriever modules.
@@
Figure 4: Success rates of ChatCFD in reproducing NACA0012 Case 1 and Nozzle Case 5 across five distinct Configurations A to E. The horizontal axis delineates these configurations, while the vertical axis quantifies the success rates. A 10-step operation indicates the CFD case successfully advanced 10 time steps or executed 10 iteration steps. Accuracy denotes precise case configuration in accordance with the source article’s specifications.
@@
Figure 4 illustrates the performance of ChatCFD across the five distinct system Configurations A through E for two benchmark test cases: NACA0012 Case 1 and Nozzle Case 5. Two primary metrics were employed for evaluation: the 10-step operational success rate, signifying the capability to complete ten simulation steps (time steps for transient cases or iterations for steady-state cases) without critical errors, and the accurate configuration success rate, indicating that the CFD case was set up in exact accordance with the specifications detailed in the source research article. The 10-step criterion was adopted to enable rapid assessment of operational stability for the system’s reflection mechanism, as target simulation durations are often unspecified in the literature. It was observed that configurations meeting this 10-step milestone were generally free of major operational faults, though minor discrepancies in boundary conditions parameters from the article’s description could still occur. For cases achieving accurate configuration, the simulation results demonstrated good agreement with published data, as detailed in Appendix D.
@@
For the NACA0012 Case 1, the data in Figure 4 show a clear progression in the 10-step operational success rate with increasing system sophistication. The baseline Configuration A achieved a 20% success rate. This rate systematically improved to 40% for Configuration B, 60% for Configuration C, and reached 80% for Configuration D, which benefited from enhanced retrieval capabilities. This trend highlights the positive impact of incrementally integrating retrieval modules and OpenFOAM-specific knowledge. However, despite its high operational success, Configuration D failed to achieve any accurate configuration success, being unable to precisely replicate the case setup from the source article. This crucial gap was addressed by Configuration E, which incorporates comprehensive article interpretation and parameter extraction as part of its advanced Stage 2 processing. With Configuration E, the accurate configuration success rate for the NACA0012 case improved significantly to 40%, while maintaining the 80% operational success rate.
The comprehensive article interpretation module in Stage 2 is integral to achieving accurate CFD configurations, particularly for cases involving non-standard boundary conditions. For instance, the NACA0012 case, as depicted in Figure 3, employs several specialized OpenFOAM boundary conditions (freestream, freestreamVelocity, and freestreamPressure) that deviate from common practices. Analysis of the LLM’s internal reasoning processes indicated that, without robust article interpretation capabilities and strict adherence to documented specifications, the model tends to default to more conventional boundary conditions. This can lead to erroneous configurations, such as using standard inlet-type conditions for inlet velocities and pressureOutlet-type conditions for outlet pressures, instead of the requisite specialized freestream conditions.
This tendency for conventional settings also extends to numerical parameters, including inlet velocity values. To mitigate these issues, the comprehensive article interpretation module was developed in Stage 2, incorporating a three-step process for analyzing articles. This process emphasizes the precise extraction of case specifications and ensures adherence to OpenFOAM’s inherent boundary condition types, as detailed in Section 2. The introduction of this module yielded substantial improvements: while the 80% success rate for the 10-step operational benchmark was maintained, the system’s capability to accurately replicate documented case configurations increased dramatically from 0% to 40%. This significant enhancement underscores the efficacy of the refined interpretation strategy in correctly identifying and implementing specialized case requirements.
In the Nozzle Case 5, as depicted in Figure 4, the introduction of the ContextRetriever module led to a substantial improvement in the 10-step operational success rate. Configurations lacking ContextRetriever (i.e., up to Configuration C) exhibited a 0% success rate for this metric. The integration of ContextRetriever in Configuration D elevated this rate to 60%. This marked improvement is attributed to ContextRetriever’s enhanced capability to diagnose and address complex coupling-related errors (refer to Appendix B for examples), a notable limitation in earlier configurations. For instance, when encountering thermophysical parameter errors in OpenFOAM (such as those related to $\Psi$, the compressibility), previous configurations often misidentified the error source as the constant/thermodynamicProperties file. However, such errors frequently originate from inconsistencies in the equation of state specifications within boundary condition entries in field files (e.g., 0/p or 0/T). The ContextRetriever module overcomes this by performing a comprehensive analysis across all relevant case directories (0/, constant/, and system/), thereby enabling more accurate error localization and effective correction strategies.
@@
The differential performance of the ContextRetriever module when applied to the NACA0012 Case 1 and the Nozzle Case 5 can be attributed to fundamental distinctions in their underlying flow physics. NACA0012 Case 1, an incompressible flow simulation employing the pisoFoam solver, exhibits a comparatively weak coupling between pressure and velocity fields. In contrast, Nozzle Case 5, a compressible flow scenario utilizing the rhoCentralFoam solver, is characterized by strong interdependencies among density, temperature, and pressure, governed by the equation of state. The advanced analytical capabilities of ContextRetriever, particularly in discerning inter-file dependencies and identifying errors related to coupled physics, proved notably more effective for the Nozzle case due to its inherent complexities. While still beneficial, the module’s impact was less pronounced for the NACA0012 case, which features a less coupled system. This observed variance underscores the proficiency of ContextRetriever in managing comprehensive physical information across a spectrum of problem complexities.
Advancing from Configuration D to E for the Nozzle Case 5 preserved the 60% 10-step operational success rate while notably enhancing the accurate configuration success rate from 0% to 30%. This pattern of improvement in configuration precision, achieved by incorporating comprehensive article interpretation in Stage 2, mirrors the observations from the NACA0012 case, indicating a consistent benefit of the complete system configuration across different flow regimes.

### v2

@@ 3 Results and discussion
The performance of ChatCFD was rigorously evaluated through a series of validation experiments encompassing three distinct categories of CFD cases: (i) 205 benchmark tutorial cases drawn from OpenFOAM tutorials and the OpenFOAM wiki, serving as standardized references for foundational validation; (ii) 110 perturbed variants, derived by systematically altering key parameters (e.g., boundary conditions, solver settings, or physical properties) in the benchmark cases to assess robustness and sensitivity; and (iii) 2 advanced literature-derived cases, directly prompted from published research articles to test real-world applicability in complex, unseen scenarios. ChatCFD demonstrated an operational success rate of 82.1% across the 315 benchmark and perturbed cases, with success defined as error-free configuration and execution leading to converged simulations. For the advanced literature-derived cases, success rates ranged from 60% to 80%, reflecting the challenges of handling intricate, domain-specific configurations without prior exposure. These results represent a substantial advancement in LLM-driven CFD automation, particularly in reducing the expertise barrier for setup and execution within the OpenFOAM framework.
3.1 Performance Evaluation on OpenFOAM Basic Cases
Figure 3: Comparison of success rates across three CFD agents (ChatCFD, MetaOpenFOAM, Foam-Agent) for 205 benchmark and 110 perturbed OpenFOAM tutorial cases. (a) Success rates by case category. (b) Distribution of test cases across categories. (c) Overall success rate comparison.
To evaluate ChatCFD, a comprehensive set of test cases was curated from OpenFOAM tutorials and the OpenFOAM wiki. Cases involving complex mesh generation techniques were filtered out, resulting in 205 benchmark tutorial cases. From these, 11 cases were selected and systematically perturbed by modifying their descriptions and adjusting parameters (e.g., boundary conditions, solver settings, or physical properties), yielding an additional 110 perturbed variant cases. This approach ensures both foundational validation and robustness testing under varied conditions. Figure 3 illustrates the performance of ChatCFD compared to MetaOpenFOAM and Foam-Agent across these cases, showing success rates by category, case distribution, and overall performance. Success is defined as the generation of error-free case files leading to converged simulations. As shown in Figure 3(c), ChatCFD achieves an overall success rate of 82.1%, significantly outperforming MetaOpenFOAM (6.2%) and Foam-Agent (42.3%). Figure 3(a) highlights ChatCFD’s consistent advantage across case categories, demonstrating its superior performance and generalization capability.
[…]
Figure 4: Performance statistics for different agents across 205 benchmark tutorial cases. (a) Average token consumption per case. (b) Distribution of token consumption. (c) Average number of reflection iterations per case. (d) Average token consumption per reflection iteration. (e) Average computational cost (in monetary terms). (f) Distribution of reflection iteration ratios, excluding zero and limit-reaching cases.
The performance of ChatCFD, MetaOpenFOAM, and Foam-Agent was evaluated through detailed metrics on 205 benchmark tutorial cases from OpenFOAM, focusing on token consumption, reflection iterations, and computational cost. Figure 4 presents a comparative analysis of these metrics, highlighting ChatCFD’s efficiency and robustness in automated CFD workflows. As shown in Figure 4(a), ChatCFD exhibits the lowest average token consumption per case, approximately half that of Foam-Agent, with MetaOpenFOAM ranking second. This efficiency underscores ChatCFD’s high success rate combined with reduced costs, making it practical for large-scale CFD applications. While ChatCFD and MetaOpenFOAM have comparable total token consumption, Figure 4(e) reveals a significant cost disparity, with ChatCFD’s computational cost being approximately 1.5 times lower than MetaOpenFOAM’s. This is primarily due to ChatCFD’s strategic use of the cost-efficient DeepSeek-V3 model (with a cost half that of DeepSeek-R1) for specific pipeline stages, as illustrated in Figure 4(b). In contrast, MetaOpenFOAM’s lower success rate necessitates frequent error corrections, generating additional file content and inflating costs. Figure 4(c) shows that MetaOpenFOAM requires the highest average number of reflection iterations per case, followed by Foam-Agent, with ChatCFD requiring the fewest. However, Figure 4(d) indicates that MetaOpenFOAM consumes fewer tokens per reflection iteration due to its simpler error-correction approach, which dilutes the high token cost of initial case setup across multiple iterations. Conversely, ChatCFD and Foam-Agent employ more sophisticated error-reflection strategies. Foam-Agent, for instance, maintains an “error correction trajectory” that logs modifications to file $i$ at iteration $j$, supplemented by contextual data from related files and reference cases. This increases its per-iteration token consumption. ChatCFD, while similarly providing contextual data, optimizes efficiency by maintaining both short-term modification trajectories and long-term reflection summaries. This dual approach enables precise filtering of irrelevant information and concise summarization of critical context, significantly reducing token consumption during LLM interactions compared to Foam-Agent, as evidenced in Figure 4(d). Figure 4(f) depicts the distribution of reflection iteration ratios, defined as the number of reflections divided by the maximum allowed iterations, reflecting each agent’s assessment of case complexity. The x-axis represents the reflection ratio, and the y-axis indicates the proportion of cases at each complexity level. Higher curves indicate stronger error-correction capabilities at a given reflection ratio. All three agents exhibit a rapid decline in solved cases followed by saturation, suggesting that increasing reflection iterations beyond a certain point yields diminishing returns. For Foam-Agent, with a reflection limit of 49 iterations, the curve flattens at approximately 50% (around 25 iterations), indicating that additional reflections beyond this threshold rarely resolve errors. ChatCFD’s curve consistently lies above the others, demonstrating more effective reflections that are likelier to correct errors. However, even for ChatCFD, reflections beyond 75% of the limit (approximately 22 iterations) often fail to resolve remaining errors, highlighting a current limitation in LLM-based CFD agents where complex case errors require advanced knowledge integration or alternative strategies beyond iterative reflection.
3.2 Evaluation and Analysis on Literature-Derived CFD Cases
Figure 5: Visualization of two literature-derived CFD cases. (a) Computational mesh for the NACA0012 airfoil case [34]. (b) Computational mesh for the Nozzle case [35]. (c) Boundary condition types for the NACA0012 case. (d) Boundary condition types for the Nozzle case.
@@ 3.2 Evaluation and Analysis on Literature-Derived CFD Cases
To evaluate ChatCFD’s ability to handle complex, real-world scenarios, two representative literature-derived cases were selected: an incompressible flow case and a compressible flow case, as illustrated in Figure 5. These cases were chosen for their comprehensive documentation in the source literature, providing detailed specifications for solver configurations, turbulence models, and mesh characteristics critical for accurate CFD reproduction. The incompressible flow case, based on Sun et al. [34], involves a NACA0012 airfoil at a 10° angle of attack, simulated using the simpleFoam solver with the Spalart-Allmaras turbulence model. The compressible flow case, drawn from Yu et al. [35], models a nozzle with a pressure ratio of 3, employing the rhoCentralFoam solver and the Spalart-Allmaras model.
Given reported challenges in modifying OpenFOAM turbulence models using large language models (LLMs), as noted by Pandey et al. [28], additional experiments were conducted to explore robustness across various turbulence models (e.g., $\epsilon, k$-$\omega$ SST). Table 1 provides a comprehensive summary of all case configurations and experimental iterations. High-fidelity computational meshes were generated to ensure accuracy. The experimental protocol was structured as follows: Cases 1 (NACA0012) and 5 (nozzle) underwent an extensive model ablation study, with 10 iterations per case across five system configurations to assess sensitivity to model variations. Cases 2 through 4, using different turbulence models, were evaluated based on the full ChatCFD system configuration, with 10 iterations each, to focus on performance under optimal settings. This rigorous methodology resulted in 130 experimental runs, forming a robust dataset for analyzing ChatCFD’s performance in complex, literature-derived scenarios. The results, detailed in subsequent analyses, demonstrate ChatCFD’s capability to adapt to sophisticated CFD configurations, complementing its strong performance on benchmark tutorial and perturbed variant cases.
@@
3.2.1 Ablation Study on ChatCFD System Configurations
@@
To systematically dissect the contributions of ChatCFD’s core components, ablation experiments were conducted across five distinct system configurations. These configurations vary in the complexity of case file generation (Stage 2, Section 2.3), the error-handling logic, and the integration of retrieval modules (ReferenceRetriever and ContextRetriever) in Stage 3 (Correcting Case Files). The configurations are defined as follows:
@@
  • Configuration A (Baseline System): Employs a simplified parameter extraction process in Stage 2, bypassing the three-step hierarchical extraction. In Stage 3, error correction is limited to a basic ’general error’ pathway without ReferenceRetriever (tutorial-based retrieval).
  • Configuration B (Baseline with ReferenceRetriever) and Simplified Setup Extraction: Retains the simplified Stage 2 extraction but enhances Stage 3 with the ReferenceRetriever module, enabling tutorial-based error correction within the general error pathway.
  • Configuration C (Enhanced Setup and Reference Retrieval): Implements the full three-step setup extraction and configuration validation in Stage 2. Stage 3 includes general error correction augmented by the ReferenceRetriever module for improved accuracy.
@@
  • Configuration D (Full Error Reflection with Simplified Extraction): Uses the simplified Stage 2 extraction but incorporates both general and persistent error correction modules in Stage 3, supported by ReferenceRetriever and ContextRetriever for comprehensive error handling.
  • Configuration E (Complete System): Represents the full ChatCFD pipeline, featuring comprehensive case setup extraction and validation in Stage 2 and complete error reflection modules in Stage 3, integrating both ReferenceRetriever and ContextRetriever for optimal performance.
@@
Figure 6: Success rates of ChatCFD across five system configurations (A–E) for literature-derived cases: NACA0012 (Case 1) and Nozzle (Case 5). The horizontal axis represents configurations, and the vertical axis quantifies success rates. The 10-step operational success rate indicates completion of 10 simulation steps (time steps for transient cases or iterations for steady-state cases) without critical errors. The accurate configuration success rate reflects precise adherence to source article specifications.
@@
Figure 6 illustrates ChatCFD’s performance across Configurations A through E for two literature-derived cases: NACA0012 (Case 1) and Nozzle (Case 5). Two metrics were evaluated: (1) the 10-step operational success rate, defined as the ability to complete 10 simulation steps (time steps for transient cases or iterations for steady-state cases) without critical errors, and (2) the accurate configuration success rate, indicating exact compliance with the solver, turbulence model, and boundary condition specifications in the source articles. The 10-step criterion enables rapid assessment of operational stability, particularly for the reflection mechanism, as literature often lacks specified simulation durations. Configurations achieving this milestone were generally free of major operational faults, though minor deviations in boundary condition parameters could persist. Cases meeting the accurate configuration criterion exhibited simulation results in strong agreement with published data, as detailed in the appendix. These results underscore ChatCFD’s robustness in handling complex, literature-derived configurations, complementing its performance on benchmark tutorial and perturbed variant cases.
@@
For the NACA0012 Case 1, the data in Figure 6 show a clear progression in the 10-step operational success rate with increasing system sophistication. The baseline Configuration A achieved a 20% success rate. This trend highlights the positive impact of incrementally integrating retrieval modules and OpenFOAM-specific knowledge. However, despite its high operational success, Configuration D failed to achieve any accurate configuration success, unable to precisely replicate the case setup from the source article. This crucial gap was addressed by Configuration E, which incorporates comprehensive article interpretation and setup extraction as part of its advanced Stage 2 processing. With Configuration E, the accurate configuration success rate for the NACA0012 case improved significantly to 40%, while maintaining the 80% operational success rate. Analysis of the DeepSeek-R1 model’s internal reasoning revealed that, without robust article interpretation and strict compliance with documented specifications in Stage 2, the LLM often defaults to generic boundary conditions and simplified parameters, such as standard inlet-type conditions for inflow velocities or pressureOutlet-type conditions for outflow pressures, instead of the specialized freestream conditions required for the NACA0012 airfoil. These erroneous configurations underscore the critical role of comprehensive case setup extraction and validation in Stage 2, which Configuration E leverages to ensure fidelity to complex CFD requirements, enhancing ChatCFD’s performance in complicated scenarios.
For the Nozzle Case 5, Figure 6 demonstrates a significant enhancement in the 10-step operational success rate with the integration of the ContextRetriever module. Configurations A through C, lacking ContextRetriever, achieved a 0% success rate for this metric, defined as completing 10 simulation steps (time steps for transient cases or iterations for steady-state cases) without critical errors. Configuration D, incorporating ContextRetriever, improved this rate to 60%. This advancement stems from ContextRetriever’s ability to diagnose complex coupling-related errors within the MCP framework, addressing a key limitation of earlier configurations. For example, errors in thermophysical parameters, such as compressibility ($\Psi$), were often misattributed to the constant/thermophysicalProperties file in prior configurations. In reality, these errors frequently arise from inconsistencies in equation of state specifications within boundary condition entries in field files (e.g., 0/p, 0/T). By analyzing all relevant directories (0/, constant/, system/), ContextRetriever enables precise error localization and effective correction, enhancing performance for complex cases.
@@
The differential impact of ContextRetriever between the NACA0012 Case 1 and Nozzle Case 5 reflects their distinct flow physics. The NACA0012 case, an incompressible flow simulation using the pisoFoam solver, exhibits weak pressure-velocity coupling, reducing the need for advanced error handling. Conversely, Nozzle Case 5, a compressible flow scenario with the rhoCentralFoam solver, involves strong interdependencies among density, temperature, and pressure, governed by the equation of state. The ContextRetriever module’s advanced analysis of inter-file dependencies proves particularly effective for the Nozzle case’s coupled physics, yielding greater performance gains compared to the less coupled NACA0012 case. This highlights ContextRetriever’s proficiency in managing complex physical interactions across diverse CFD scenarios.
Advancing to Configuration E, which integrates comprehensive article interpretation in Stage 2, maintained the 60% 10-step operational success rate for Nozzle Case 5 while increasing the accurate configuration success rate from 0% to 30%. This metric reflects precise adherence to source specifications. Consistent with findings for the NACA0012 case, Configuration E’s enhanced setup extraction and validation ensure greater configuration fidelity across flow regimes, reinforcing ChatCFD’s robustness in literature-derived cases.
Figure 7: ChatCFD’s performance metrics for accurate configuration of literature-derived NACA0012 Case 1 and Nozzle Case 5. Metrics include: number of reflection iterations (#reflection), total token consumption for DeepSeek-R1 and DeepSeek-V3 models (R1/V3 total token), execution cost (Cost \$), and average token consumption per reflection round for both models (R1/V3 token/#reflection).
The performance disparity between the literature-derived NACA0012 Case 1 and Nozzle Case 5 arises from two key factors: (1) the increased complexity of managing numerous files with intricate content in compressible flows, and (2) the stringent requirement for consistency across coupled thermophysical models (e.g., 0/p, 0/T, constant/thermophysicalProperties). The Nozzle case, a compressible flow simulation using rhoCentralFoam, involves strong coupling among pressure, density, temperature, and the equation of state, necessitating precise dimensional and physical alignment across governing equations and material properties. This contrasts with the NACA0012 case, an incompressible flow simulation using pisoFoam with weaker pressure-velocity coupling, which imposes fewer constraints. Additionally, an implicit bias in LLM training data, predominantly focused on incompressible flow scenarios, leads to erroneous defaults in compressible cases. For instance, in OpenFOAM, incompressible flows use kinematic pressure ($p_{k}$, dimensions $L^{2}T^{-2}$)—dynamic pressure divided by density—while compressible flows require absolute thermodynamic pressure ($p$, dimensions $ML^{-1}T^{-2}$). The DeepSeek-R1 model’s tendency to apply incompressible conventions to compressible cases introduces dimensional errors, such as incorrect pressure settings, requiring additional reflection iterations for correction.
Figure 7 quantifies these challenges: the Nozzle case requires over twice the reflection iterations, triple the DeepSeek-R1 tokens, and six times the DeepSeek-V3 tokens compared to the NACA0012 case, resulting in approximately triple the execution cost. The disproportionate DeepSeek-V3 token increase (6x vs. 3x for DeepSeek-R1) reflects frequent dimensional inconsistencies and persistent errors, necessitating more invocations of the costlier DeepSeek-V3 model. Figure 7(c) confirms this, showing a DeepSeek-V3 token multiplier of  2.7x per reflection for the Nozzle case, compared to  2x for DeepSeek-R1, relative to the NACA0012 case. These metrics highlight the computational burden of complex flow simulations and underscore ChatCFD’s ability through advanced error correction.
3.2.2 Impact of Turbulence Models on ChatCFD Performance
Figure 8 presents ChatCFD’s performance across four turbulence models—Spalart-Allmaras (SA), $\epsilon, k$-$\omega$ SST, and RNG k-$\epsilon$—applied to the literature-derived NACA0012 cases (Cases 1 to 4). The analysis focuses on simulations achieving accurate configuration per published specifications, excluding those with only basic operational success (e.g., 10-step operational success rate). All results were obtained using the complete ChatCFD system (Configuration E).
As shown in Figure 8(a), success rates for accurate configuration vary significantly: SA at 40%, k-$\omega$ SST at 30%, k-$\epsilon$ at 20%, and RNG k-$\epsilon$ at 10%. This demonstrates ChatCFD’s ability to handle diverse turbulence models, addressing LLM challenges noted in OpenFOAMGPT [28] through structured knowledge integration and RAG. Performance metrics in Figure 8(a,b) reveal consistent profiles for SA, k-$\epsilon$, and k-$\omega$ SST, with an average of  6 reflection iterations, comparable DeepSeek-R1 token usage, and costs. In contrast, RNG k-$\epsilon$ requires  23.5 reflections—nearly 4x higher—elevating token consumption and costs.
Figure 8(c) attributes these disparities to configuration file token counts and model prevalence in OpenFOAM tutorials. While file sizes differ modestly ( 1,800 tokens for SA vs.  2,000 for others), tutorial distribution is more influential: SA, k-$\epsilon$, and k-$\omega$ SST appear in 28+ tutorials each, versus one for RNG k-$\epsilon$. RNG k-$\epsilon$’s performance deficit stems from: (1) its rarity in engineering applications, leading to sparse LLM training data and poor zero-shot configuration; and (2) limited ReferenceRetriever efficacy due to few OpenFOAM tutorials, hindering guidance for parameters in constant/turbulenceProperties, system/fvSchemes, and system/fvOption. This necessitates greater reliance on the LLM’s limited intrinsic knowledge, reducing performance relative to common models.
Figure 8: ChatCFD’s average performance for accurate configuration of NACA0012 Cases 1 to 4 across turbulence models (Spalart-Allmaras, $\epsilon, k$-$\omega$ SST, RNG k-$\epsilon$). Metrics shown include: accuracy (success rate of precise configuration according to literature), number of reflection iterations (#reflection), total token consumption of the DeepSeek-R1 model (R1 total token), execution cost (Cost \$), token count of configuration files (#token of config files), and the number of cases incorporating each turbulence model in OpenFOAM tutorials (#Case of model in OF tutorial).
3.3 Future Potential of ChatCFD in Multi-Agent CFD Workflows
Building on the insights from ChatCFD’s performance in benchmark tutorial, perturbed variant, and literature-derived cases, ongoing work explores its potential as a foundational component in advanced engineering workflows. Beyond functioning as a standalone tool for CFD automation, ChatCFD is being adapted through the Model Context Protocol (MCP), a standardized framework for integrating specialized AI agents into a cohesive multi-agent system. MCP facilitates seamless communication and data exchange among agents by defining structured protocols for context sharing, enabling ChatCFD to collaborate with other intelligent tools to address complex, interdisciplinary engineering tasks.
As illustrated in Figure 9, this multi-agent workflow enables an end-to-end design cycle for CFD applications. The process begins with a 3D modeling agent that generates a three-dimensional geometry from a natural language description or a simple image, such as a sketch or photograph. A subsequent meshing agent automatically discretizes the geometry into a high-fidelity computational mesh, leveraging tools like commercial mesh generators or open-source alternatives (e.g., blockMesh, snappyHexMesh). This mesh is then passed to the ChatCFD agent, which autonomously configures solver parameters, boundary conditions, and turbulence models (e.g., simpleFoam with Spalart-Allmaras or rhoCentralFoam with k-$\omega$ SST) to execute the CFD simulation, computing key performance metrics such as drag coefficients or pressure distributions. Finally, an optimization agent analyzes the simulation results, providing insights for iterative design refinement, such as adjusting geometry or boundary conditions to minimize drag or enhance flow stability. This MCP agent is available to use and can be found online (https://www.bohrium.com/apps/designagent0812).
This integrated approach highlights ChatCFD’s transformative potential beyond simulation automation. By serving as a standardized computational module within a collaborative multi-agent ecosystem, ChatCFD can contribute to broader engineering design processes, from conceptual design to performance optimization. The MCP framework ensures interoperability, allowing ChatCFD to adapt to diverse tasks, such as aerodynamics, heat transfer, or multiphase flow simulations, by interfacing with agents specialized in geometry generation, mesh refinement, or post-processing. This vision positions ChatCFD as a cornerstone for future AI-driven engineering workflows, paving the way for scalable, interdisciplinary applications in scientific discovery and industrial innovation.
Figure 9: ChatCFD integrated within a multi-agent workflow via the Model Context Protocol (MCP), enabling transformation from a single sentence or image to a complete CFD simulation.
4 Conclusion
This paper introduces ChatCFD, an innovative LLM-based CFD agent system for automating OpenFOAM simulations. By incorporating a multi-agent architecture, domain-specific knowledge bases, and advanced error reflection mechanisms, ChatCFD achieves end-to-end automation for reproducing literature results, including both incompressible (NACA0012 airfoil) and compressible (supersonic nozzle) flows. The simulations accurately reproduce velocity contours, pressure coefficients, and schlieren images, aligning closely with experimental data. ChatCFD’s innovations include: (1) an interactive multi-modal input mechanism integrating PDFs, meshes, and dialogue with physical constraints, providing a ChatGPT-like interface for guided literature analysis, case selection, and parameter specification, reducing operational barriers; (2) specialized AI agents leveraging CFD expertise for task interpretation, feature library construction via OpenFOAM preprocessing, and iterative reflection-refinement using DeepSeek-R1 for reasoning and DeepSeek-V3 for precise modifications; and (3) a robust error reflection system with ReferenceRetriever (tutorial templates) and ContextRetriever (case configurations) for real-time error classification and resolution.
Experimental validations on 315 basic cases (205 benchmark tutorials, 110 perturbed variants) yield an 82.1% operational success rate, outperforming MetaOpenFOAM (6.2%) and Foam-Agent (42.3%), with lower success in multiphase/combustion due to interdisciplinary complexity. On literature-derived cases, ChatCFD achieves 60–80% success, with ablation studies showing comprehensive interpretation boosts accurate configuration from 0% to 40% for NACA0012 and 30% for Nozzle, while ContextRetriever improves operational success from 0% to 60% in Nozzle by resolving coupled-physics errors. Turbulence model analyses reveal 40% success for common models (e.g., Spalart-Allmaras) vs. 10% for rare ones (e.g., RNG k-$\epsilon$), attributed to data scarcity. Physics coupling studies indicate compressible flows demand tripled resources due to interdependencies, and LLM bias toward incompressible setups causes dimensional errors in complex cases. These findings highlight ChatCFD’s robustness, efficiency, and cost-effectiveness, establishing benchmarks for AI-driven CFD.
ChatCFD effectively addresses the critical gaps in CFD automation agents. First, it integrates domain-specific knowledge through structured RAG and expert-designed agents, mitigating LLM training scarcity to enable precise boundary and initial condition specifications. Second, its modular, MCP-compatible design supports collaborative multi-agent networks, integrating specialized meshing agents to handle complex geometries beyond blockMesh. Third, by leveraging extensive literature-derived corpora with detailed specifications and benchmarks, it will facilitate scalable testing and iterative agent evolution, harnessing millions of CFD papers for refinement and validation in the future. These advancements accelerate scientific discovery in fluid mechanics and engineering by automating hypothesis testing, parameter optimization, and multiphysics simulations.
ACKNOWLEDGEMENTS
This work is sponsored by the National Natural Science Foundation of China, Grant No. 92470127, and the Overseas Postdoctoral Talents Program in Guangdong Province.
AUTHOR DECLARATIONS
Conflict of Interest
The authors declare no conflicts of interest.
Author Contributions
E Fan: Methodology, Software, Data curation, Formal analysis, Resources, Writing - original draft. Kang Hu: Methodology, Software, Data curation, Formal analysis, Resources. Zhuowen Wu: Methodology. Jiangyang Ge: Data curation, Formal analysis. Jiawei Miao: Software, Resources. Yuzhi Zhang: Conceptualization, Software, Resources. He Sun: Resources, Writing - Review. Weizong Wang: Resources, Writing - Review, Funding acquisition. Tianhan Zhang: Conceptualization, Methodology, Formal analysis, Resources, Writing - Review & Editing, Visualization, Supervision, Funding acquisition
Data Availability
The data supporting the findings are available at https://github.com/ConMoo/ChatCFD.
Appendix A Appendix: detailed analysis of two literature-derived cases